# Boosting Adverse Weather Crowd Counting via Multi-queue Contrastive Learning

### v2

@@ Abstract
Currently, most crowd counting methods have outstanding performance under normal weather conditions. However, they often struggle to maintain their performance in extreme and adverse weather conditions due to significant differences in the domain and a lack of adverse weather images for training. To address this issue and enhance the model’s robustness in adverse weather, we propose a two-stage crowd counting method. Specifically, in the first stage, we introduce a multi-queue MoCo contrastive learning strategy to tackle the problem of weather class imbalance. This strategy facilitates the learning of weather-aware representations by the model. In the second stage, we propose to refine the representations under the guidance of contrastive learning, enabling the conversion of the weather-aware representations to the normal weather domain. While significantly improving the robustness, our method only marginally increases the weight of the model. In addition, we also create a new synthetic adverse weather dataset. Extensive experimental results show that our method achieves competitive performance.
keywords:
crowd counting , adverse weather , contrastive learning.
journal: Pattern Recognition
@@
[a1]organization=School of Computer Science and Engineering, Nanjing University of Science and Technology,city=Nanjing, postcode=210094, state=Jiangsu, country=China
[…]
[a2]organization=Guangxi Key Lab of Multi-source Information Mining & Security, Guangxi Normal University,city=Guilin, postcode=541004, state=Guangxi, country=China
@@ 1 Introduction
Crowd counting has attracted much attention recently due to its wide range of applications such as public safety, video surveillance, and traffic control. Currently, most of the crowd counting methods [45; 19; 22] are able to estimate the number of crowds well on the images recorded under normal weather conditions. However, when it comes to adverse weather conditions such as rain, haze, and snow, these methods face challenges in maintaining their performance due to domain differences and the limited availability of adverse weather images. What’s worse, labeled adverse weather crowd images are exceedingly rare. For example, in the JHU-Crowd++ dataset [34], the number of images under rain conditions accounts for only 3% of the total dataset.
@@
To mitigate the negative influence of adverse weather, an intuitive remedy is to pre-process the images using image restoration modules before counting. Unfortunately, even when restoration modules can mitigate the obscuration caused by adverse weather, the restored images still exhibit significant domain differences from normal weather images. Moreover, the additionally introduced classification and enhancement modules significantly increase the computational burden. Training a separate model for each weather category is also not feasible because the limited samples of adverse weather conditions are insufficient to support the training. Huang et al. [10] introduces a transformer-based approach that addresses the issue of generating complementary information from image-specific degradation without the need for customized image enhancement modules. However, this transformer-based approach not only poses a heavy computational burden but also ignores the issue of weather class imbalance in the dataset.
@@
The aim of this study is to enhance the robustness of crowd counting models in unknown adverse weather conditions while maintaining their performance under normal weather conditions without significantly increasing model complexity. To achieve this objective, we approach the problem of crowd counting, which includes both normal and adverse weather conditions, as an imbalanced multi-domain learning task. The key prerequisite for a model to tackle multi-domain learning is to have perception capabilities across different domains, meaning that the extracted information from each domain should possess discriminative characteristics. This aligns with the objective of contrastive learning [3; 8; 30]. Therefore, in this paper, we propose a two-stage method called Multi-queue Contrastive Learning (MQCL). This approach enables the backbone model to directly extract weather-aware representations, which are further refined by a refiner module. In the first stage, we use contrastive learning to distinguish the characteristics of different weather. However, since the class imbalance mentioned above, the gradient of the loss function of the vanilla contrastive method would be dominated by normal weather images, resulting in poor performance of representations and difficulty in refining and counting. To tackle such an imbalance problem, we design a simple yet effective contrastive learning method called multi-queue MoCo, which replaces the standard single queue in MoCo [8] with multiple queues, providing class-balanced key vectors. In the second stage, all normal weather samples in the queues are treated as positive samples of contrastive learning. Thus the refiner can be guided to convert the representations of adverse weather images to the domain of normal images. Benefiting from the effective representation learning, the refiner and the decoder can be designed to be light to meet our target of not significantly increasing the weight. Comparison of weight and accuracy of MQCL and state-of-the-art methods is shown in Fig. 1(a/b).
Figure 1: Trade-off of model weight and accuracy between our proposed method and state-of-the-art methods under adverse weather (a) and normal weather (b) conditions on the JHU-Crowd++ dataset. The radius of the circle is proportional to the number of parameters of the model.
In addition, as currently only one publicly available real-world dataset, JHU-Crowd++ [34], contains adverse weather images and annotations, we synthesized a new dataset called NWPU-Weather based on the NWPU-Crowd dataset [42] with rainy and hazy scenes. This dataset aims to facilitate research on crowd counting in adverse weather conditions. Several representative counting networks are benchmarked to provide an overview of the state-of-the-art performance. Codes and the NWPU-Weather dataset are available at: https://anonymous.4open.science/r/MQCL-B46E/.
@@
  • 1.
    To boost the robustness of the model under adverse weather conditions while maintaining normal-weather performance, we propose a lightweight two-stage method, achieving significant improvement compared to the baseline.
  • 2.
@@
  • 4.
    To provide the crowd counting field with more experimental samples in adverse weather, we synthesize a new adverse weather crowd counting dataset called NWPU-Weather. Extensive experimental results show that our method achieves competitive results.
@@ 2.1 Crowd Counting under Normal Conditions
On the other hand, the target of localization-based methods is to predict head positions. Song et al. [35] propose a model called P2PNet to directly predict a set of point proposals to represent heads in an image, being consistent with the human annotation results. They also proposed a new metric called density Normalized Average Precision (nAP) to provide more comprehensive and more precise performance evaluation. CLTR [21] is a concurrent work with P2PNet, introducing the Transformer architecture to the field of crowd localization and proposing a KMO-based Hungarian matcher to reduce the ambiguous points and generate more reasonable matching results.
@@ 2.2 Crowd Counting under Adverse Conditions
Existing deep-learning-based methods have achieved unprecedented success with crowd counting, but their performance degraded severely under adverse conditions (e.g., adverse weather) due to the disturbance to the brightness and gradient consistency. However, few research efforts have been made into this problem. Additional class conditioning blocks are utilized by [34] to augment the backbone module, which is trained via cross-entropy error using labels available in the dataset. AWCC [10] enables the model to extract weather information according to the degradation via learning adaptive query vectors, but the weight of the model is significantly increased due to the introduction of a Transformer-based module. Kong et al. [15] proposes a single-stage hazy-weather crowd counting method based on direction-aware attention aggregation. However, their method only focuses on the performance in hazy scenes and cannot handle various unknown weather conditions.
@@ 2.3 Contrastive Learning
Contrastive learning [30; 3] has attracted much attention due to its success in unsupervised representation learning. The target of it is to maximize the similarity of the representations between positive pairs while minimizing that of negative pairs. He et al. [8] builds a dynamic dictionary with a queue and a moving-averaged encoder to enable large-scale contrastive learning with dramatically low demand for memory. [14] investigates the contrastive loss and adapted contrastive learning to the field of supervised learning.
@@ 3 Proposed Method
3.1 Framework Overview
The architecture of our method is illustrated in Fig. 2. As discussed earlier, we aim to directly enable the crowd counting backbone model to learn weather-aware feature representations and then refine it with a light-weight refiner. Thus, the decoder can focus on a single domain and generate high-quality density map, the sum of which is the estimated number. Since the refiner is premised on stable and consistent representation while it keeps evolving and is not stable during the representation learning stage, we divide the training into two stages to separate these two targets, namely Weather-aware Representation Learning (WRL) stage and Contrastive-learning-guided Representation Refining (CRR) stage, respectively. After training is completed, the model of the CRR stage is used for inference.
In the WRL stage, contrastive learning is utilized to enable the encoder to learn weather-aware representations. The contrastive learning strategy of the WRL stage is based on our proposed multi-queue MoCo, which can tackle the problem of weather class imbalance. The weights of the encoder, the decoder and the queues obtained during the WRL stage will be retained for the CRR stage. In the CRR stage, the encoder is freezed and the main task of this stage is to train a refiner. By treating the normal weather samples as positive examples of contrastive learning, the refiner is guided to pull the adverse weather representations towards normal weather domain. Finally, high-quality density map can be generated to realize precise counting. Note that the tiny version of ConvNeXt [26] is used as the encoder in our work. In terms of the decoder’s structure, we remain consistent with MFA [23], which primarily consists of two transposed convolutional layers.
@@ 3.2 Weather-aware Representation Learning
The model architecture of the WRL stage is illustrated in the upper half of Fig. 2. The contrastive learning method is utilized to train the encoders and endow it with the capability to extract weather-aware representations, i.e., images with similar weather conditions correspond to similar representations, whereas those with dissimilar conditions correspond to more distant representations.
Our training structure is the same as MoCo [8], which consists of an encoder $\rm E_{Q}$, a momentum-updated encoder $\rm E_{K}$ and a decoder. The encoders $\rm E_{Q}$ and $\rm E_{K}$ extracts representations $\rm R_{Q}$ and $\rm R_{K}$, respectively, from different augmentations $\rm I_{Q}$ and $\rm I_{K}$ of an image. The representations $\rm R_{Q}$ and $\rm R_{K}$ will be further processed by our proposed multi-queue MoCo strategy, which tackles the problem of weather class imbalance and grant the representations with weather-aware nature, to obtain the contrastive loss. The details of multi-queue MoCo will be elaborated in Section 3.3. To ensure that the representation simultaneously contains crowd information, the target of the decoder is set as generating a density map under the supervision of the Bayesian loss [28] according to the representation $\rm R_{Q}$. The total loss of the WRL stage is:
@@ 3.3 Multi-queue MoCo
To tackle the problem, we propose multi-queue MoCo, the architecture of which is illustrated in Fig. 3. Similar to most of the contrastive methods, each image undergoes data augmentation multiple times, and the representations originating from the same image as the anchor are treated as positive samples while those from different images are treated as negative samples. This strategy may lead to a situation where samples with same weather conditions from different images are treated as negatives. This seems to be inconsistent with our motivation. But fortunately, research conducted by [41] indicates that contrastive learning has the tolerance to semantically similar negative samples. In light of this, we conducted extensive experiments and the results indicate that such a strategy outperforms the positive/negative partitioning strategy based on weather labels.
@@ 3.4 Contrastive-learning-guided Representation Refining
The multiple queues obtained during the WRL stage have stored a sufficient number of vectors from various weather conditions. In this paper, we propose to preserve and freeze these queues and employ contrastive learning to guide the refiner to convert the representation. In this stage, all the vectors in the normal-weather sub queue are treated as positive samples and those in other sub queues are treated as negatives. Since there is no longer a need to generate vectors K, the encoder $\rm E_{K}$ and projection head $\rm P_{K}$ are discarded. In order to maintain stable representations, the parameters of encoder $\rm E_{Q}$ are fixed at this stage. The projection head $\rm P_{Q}$ is preserved, fixed, and moved behind the refiner. We denote the vector projected from the refined representation as $\rm Q^{\prime}$. The model structure of the CRR stage is shown in the lower half of Fig. 2. The loss function of contrastive learning in this stage is calculated as follows:
@@ 4.1 Experiment Setup
The ConvNeXt [26] is employed as the encoder and we use the tiny version for simplicity. The structure of ConvNeXt-T is: C = (96, 192, 384, 768), B = (3, 3, 9, 3), where C stands for the number of channels and B is the number of ConvNeXt blocks in each stage. The decoder’s structure is consistent with MFA [23], which is an upsampling module composed of two transposed convolutional layers. The encoder $\rm E_{Q}$ and encoder $\rm E_{K}$ both have the same structure as ConvNeXt-T (the classification head is discarded) and the structure of the decoder keeps consistent with MFA [23], which primarily consists of two transposed convolutional layers. The pre-trained weights of ImageNet-22K [31] are loaded as the initial parameters. Two-layer MLPs are employed in the projection heads, the output dimension of which is 2048 and 128. The refiner consists of three ConvNeXt blocks with input and output dimensions of 768, initialized with random parameters.
@@
As mentioned above, each image in a batch undergoes augmentation twice. Specifically, we random crop the image with a size of 256 $\times$ 256, and horizontal flipping is applied for a probability of 50%. In the multi-queue structure, the number of the sub queues is equal to the number of weather classes, and the length of each sub queue is set to 1024. The AdamW optimizer [27] is adopted both in the WRL and CRR stage, the learning rate is scheduled by a cosine annealing strategy and the initial learning rate is $10^{-4}$. The weight decay is set to $10^{-3}$ and the batch size is $16$. $\lambda_{1}$ and $\lambda_{2}$ in the loss function are both set to 10 and the temperature $\tau$ is set to 0.05.
@@
Following the convention of existing works [19; 22], we adopt Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) as the metrics to evaluate the methods.
Figure 4: Examples of normal (a), hazy (b) and rainy (c) weather images in the NWPU-Weather dataset.
Table 1: The distribution of the number of images under different weather conditions in the NWPU-Weather dataset.
Stage Normal Haze Rain
Train 2365 120 124
Val 446 31 23
Test 442 28 30
Total 3253 179 177
4.2 NWPU-Weather Dataset
Considering the current scarcity of crowd counting datasets containing adverse weather scenarios and labels, we synthesize an adverse weather crowd counting dataset containing hazy and rainy scenes based on the NWPU-Crowd dataset [42], namely NWPU-Weather. Since the test set of the NWPU-Crowd dataset is not publicly available, we extract part of the original training set as the test set. The first 1-2609 images from the original dataset are used as the training set, the images from 2610 to 3109 are used as the test set, and the validation set still consists of images from 3110 to 3609.
Table 2: Quantitative results comparing with the state-of-the-art methods on the NWPU-Weather dataset. The numbers in parentheses represent the rankings of our method.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
CSRNet [19] (CVPR 18) 74.8 200.7 175.3 730.2
BL [28] (ICCV 19) 69.9 248.1 137.1 339.4
DM-Count [40] (NeurIPS 20) 80.6 319.1 153.6 338.2
KDMG [39] (PAMI 20) 108.0 318.6 151.9 328.4
MAN [22] (CVPR 22) 64.1 259.1 105.9 264.1
LDFNet [5] (TIP 24) 65.2 266.2 114.2 298.4
ConvNeXt [26] (CVPR 22) 69.3 264.0 108.2 286.3
MQCL (Ours) 61.6 (1) 245.5 (2) 93.3 (1) 250.1 (1)
Consistent with the JHU-Crowd++ dataset, our NWPU-Weather dataset keeps the imbalance of weather types, the hazy and rainy scenes are synthesized by a probability of 5%, respectively. The distribution of the weather conditions in the dataset is shown in Table 1. We follow the approach outlined in [17] to synthesize the weather scenarios. The intensity, density, and angle of the rain are set as random values. The depth maps required during the haze synthesis process are estimated by ZoeDepth [2] and the intensity of the haze is also random. Examples of the images in the NWPU-Weather images are shown in the Fig. 4.
Except for the MQCL, we also benchmark several representative counting networks, providing an overview of the state-of-the-art performance. Note that since the model structure of AWCC-Net [10] is not fully elucidated and the code for training is not available, we do not provide the performance. The structure our baseline ConvNeXt in Table 3 is consistent of the original model in the work [26], and we only replace the classification head with the same decoder used in MQCL. The quantitative results of counting accuracy are listed in Table 2. Compared with the baseline ConvNeXt, MQCL exhibits a significant performance improvement under adverse weather conditions. MAE and RMSE are decreased by 13.8% and 12.6%, suggesting that the proposed representation learning and refining strategy do indeed boost the robustness of the model under adverse weather conditions. Moreover, MQCL can also improve the performance under normal conditions, decreasing MAE and RMSE by 11.1% and 7.0%, respectively. We believe that this can be attributed to the contrastive learning strategy employed in this paper, which considers augmentations from the same image as positive examples. This strategy not only assists the model in weather perception but also strengthens the model’s ability to recognize different scenes. Sampling from the same image ensures that the positive examples not only share the same weather conditions but also possess similar scene characteristics. Compared with the previously best method MAN, MQCL has also achieved significant performance improvement. The MAE/RMSE under normal and adverse weather conditions have reduced by 3.9%/5.2% and 11.9%/5.3%, respectively.
@@ 4.2 NWPU-Weather Dataset
Table 3: Quantitative results comparing with the state-of-the-art methods on the JHU-Crowd++ dataset. The numbers in parentheses represent the rankings of our method.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
SFCN [43] (CVPR 19) 71.4 225.3 122.8 606.3
BL [28] (ICCV 19) 66.2 200.6 140.1 675.7
LSCCNN [32] (PAMI 20) 103.8 399.2 178.0 744.3
CG-DRCN-V [34] (PAMI 20) 74.7 253.4 138.6 654.0
CG-DRCN-R [34] (PAMI 20) 64.4 205.9 120.0 580.8
UOT [29] (AAAI 21) 53.1 148.2 114.9 610.7
GL [38] (CVPR 21) 54.2 159.8 115.9 602.1
CLTR [21] (ECCV 22) 52.7 148.1 109.5 568.5
MAN [22] (CVPR 22) 46.5 137.9 105.3 478.4
AWCC-Net [10] (ICCV 23) 47.6 153.9 87.3 430.1
LDFNet [5] (TIP 24) 52.8 156.5 112.4 530.1
ConvNeXt [26] (CVPR 22) 52.7 154.9 105.1 561.4
MQCL (Ours) 50.5 (3) 152.0 (4) 96.5 (2) 522.8 (3)

### v3

@@ Abstract
Currently, most crowd counting methods have outstanding performance under normal weather conditions. However, our experimental validation reveals two key obstacles limiting the accuracy improvement of crowd counting models: 1) the domain gap bewteen the adverse weather and the normal weather images; 2) the weather class imbalance in the training set. To address the problems, we propose a two-stage crowd counting method named Multi-queue Contrastive Learning (MQCL). Specifically, in the first stage, our target is to equip the backbone network with weather-awareness capabilities. In this process, a contrastive learning method named multi-queue MoCo designed by us is employed to enable representation learning under weather class imbalance. After the first stage is completed, the backbone model is “mature” enough to extract weather-related representations. On this basis, we proceed to the second stage, in which we propose to refine the representations under the guidance of contrastive learning, enabling the conversion of the weather-aware representations to the normal weather domain. Through such representation and conversion, the model achieves robust counting performance under both normal and adverse weather conditions. Extensive experimental results show that, compared to the baseline, MQCL reduces the counting error under adverse weather conditions by 22%, while introducing only about 13% increase in computational burdon, which achieves state-of-the-art performance.
@@ 1 Introduction
Crowd counting has attracted much attention due to its wide range of applications such as public safety, video surveillance, and traffic control. Currently, most of the crowd counting methods [39, 17, 19] are able to estimate the number of crowds well on the images recorded under normal weather conditions. However, adverse weather such as haze, rain and snow can alter the texture and edge features of objects and even affect the clothing of people. The domain gap frequently results in counting difficulties. The limited robustness to adverse weather fundamentally bottlenecks the advancement of crowd counting methods. What’s worse, labeled adverse weather crowd images are exceedingly rare. For example, as shown in the pie charts in Figure 1 (a), in the JHU-Crowd++ dataset [30], the number of images under rain conditions accounts for only 3% of the total dataset. Such imbalance can severely damage model performance. As shown in Figure 1 (c) and (d), when training data imbalance increases, existing crowd counting models, such as MAN [19], exhibit only negligible performance improvement on normal-weather images but severe degradation on adverse-weather images.
Figure 1: (a) and (b) shows the the weather condition distribution pie charts of the JHU-Crowd++ dataset and the GCC dataset, respectively. (c) and (d) present the counting accuracy of different models on normal-weather and adverse-weather images respectively, under varying degrees of weather-class imbalance in the training data. The horizontal axis represents the number of normal/hazy/rainy/snowy weather images in the training set. The vertical axis represents the Mean Absolute Error (MAE) of models on the JHU-Crowd++ test set. All training data are sampled from the JHU-Crowd++ training set.
@@
To mitigate the negative influence of adverse weather, an intuitive remedy is to pre-process the images using image restoration modules before counting. Unfortunately, even when restoration modules can mitigate the obscuration caused by adverse weather, the restored images still exhibit significant domain differences from normal weather images. For example, the restoration module cannot eliminate the occlusion caused by objects such as raincoats or umbrellas. Moreover, the additionally introduced classification and enhancement modules significantly increase the computational burden. Training a separate model for each weather category is also not feasible because the limited samples of adverse weather conditions are insufficient to support the training. AWCC-Net [10] is currently the only work that dives into how to improve model performance under diverse adverse weather conditions. It introduces a transformer-based approach that addresses the issue of generating complementary information from image-specific degradation without the need for customized image enhancement modules. However, its primary limitation is the disregard for performance degradation under weather-class imbalance. As shown in Figure 1 (d), it achieves significant performance improvement in adverse weather conditions, but still suffers severe degradation when class imbalance intensifies. Moreover, the Transformer-based design poses a heavy computational burden to the model.
@@
We aim to design a method that satisfies three objectives: (i) achieve high accuracy in adverse weather conditions even when training data is class-imbalanced; (ii) maintain high accuracy in normal weather conditions; (iii) be light-weight. To achieve above objectives, we approach the problem of crowd counting, which includes both normal and adverse weather conditions, as an imbalanced multi-domain learning task. The key prerequisite for a model to tackle multi-domain learning is to have perception capabilities across different domains, meaning that the extracted information from each domain should possess discriminative characteristics. This aligns with the objective of contrastive learning [3, 8, 27]. In this paper, we propose a two-stage method called Multi-queue Contrastive Learning (MQCL). This approach enables the backbone model to directly extract weather-aware representations, which are further refined by a refiner module. In the first stage, we use contrastive learning to distinguish the characteristics of different weather. However, since the class imbalance mentioned above, normal weather samples dominate the key vector distribution compared to adverse weather samples during training, resulting in poor performance of representations and difficulty in refining and counting. To tackle such problem, we design a simple yet effective contrastive learning method called multi-queue MoCo, which replaces the standard single queue in MoCo [8] with multiple queues, providing class-balanced key vectors. In the second stage, all normal weather samples in the queues are treated as positive samples of contrastive learning. Thus the refiner can be guided to convert the representations of adverse weather images to the domain of normal images.
The following example can more concretely explain how MQCL works: Owing to the training during the WRL stage, the backbone model can encode humans with clear textures under normal weather as representation $A$, and humans wearing raincoats in rainy weather as representation $B$. Furthermore, due to the training in the CRR stage, the refiner learns that within the context of crowd counting, representation $B$ is essentially equivalent to representation $A$, as both denote a single individual. It thus transforms representation $B$ into $A$. Therefore, even under rainy conditions, the crowd counting head can still provide accurate count predictions.
As a plug-and-play method, MQCL perfectly satisfies the three aforementioned objectives: achieving high accuracy in both normal and adverse weather conditions, while demonstrating broad adaptability to diverse backbone architectures and high computational efficiency as only a light-weight refiner is introduced. As shown in Figure 1 (c) and (d), MQCL achieves strong performance not only under balanced data conditions, but also demonstrates significantly greater robustness than other methods when faced with increased data imbalance.
@@
  • To boost the robustness of the model under adverse weather conditions while maintaining normal-weather performance, we propose a light-weight two-stage method called MQCL, achieving significant improvement compared to the baseline.
@@
  • Extensive experimental results show that our method achieves competitive results.
@@ 2.1 Crowd Counting under Normal Conditions
On the other hand, the target of localization-based methods is to predict head positions. Song et al. [31] propose a model called P2PNet to directly predict a set of point proposals to represent heads in an image, being consistent with the human annotation results. They also proposed a new metric called density Normalized Average Precision (nAP) to provide more comprehensive and more precise performance evaluation. CLTR [18] is a concurrent work with P2PNet, introducing the Transformer architecture to the field of crowd localization and proposing a KMO-based Hungarian matcher to reduce the ambiguous points and generate more reasonable matching results. APGCC [2] provides clear and effective guidance for proposal selection and optimization, addressing the core issue of matching uncertainty in point-based crowd counting.
@@ 2.2 Crowd Counting under Adverse Conditions
Existing deep-learning-based methods have achieved unprecedented success with crowd counting, but their performance degraded severely under adverse conditions due to the disturbance to the brightness and gradient consistency. Additional class conditioning blocks are utilized by [30] to augment the backbone module, which is trained via cross-entropy error using labels available in the dataset. AWCC [10] enables the model to extract weather information according to the degradation via learning adaptive query vectors, but the weight of the model is significantly increased due to the introduction of a Transformer-based module. Kong et al. [14] proposes a single-stage hazy-weather crowd counting method based on direction-aware attention aggregation. RestoCrowd [15] is also a method focusing on haze-weather crowd counting, which leverages the image restoration coadjutant learning to achieve accurate counting without weather-label. However, these methods exhibit significant limitations, such as the inability to handle multiple types of unknown weather conditions or the neglect of class imbalance in weather categories within the dataset. In this paper, MQCL not only be able to handle various types of unknown weather conditions, but also explicitly addresses the issue of weather category imbalance as one of its core design considerations.
@@ 2.3 Contrastive Learning
Contrastive learning [27, 3] has attracted much attention due to its success in unsupervised representation learning. The target of it is to maximize the similarity of the representations between positive pairs while minimizing that of negative pairs. He et al. [8] builds a dynamic dictionary with a queue and a moving-averaged encoder to enable large-scale contrastive learning with dramatically low demand for memory. [13] investigates the contrastive loss and adapted contrastive learning to the field of supervised learning.
@@ 3 Proposed Method
3.1 Framework Overview
The architecture of our method is illustrated in Figure 2. As discussed earlier, we aim to directly enable the crowd counting backbone model to learn weather-aware feature representations and then refine it with a light-weight refiner. Since the refiner is premised on stable and consistent representation while it keeps evolving and is not stable during the representation learning stage, we divide the training into two stages to separate these two targets, namely Weather-aware Representation Learning (WRL) stage and Contrastive-learning-guided Representation Refining (CRR) stage, respectively. After training is completed, the model of the CRR stage is used for inference.
In the WRL stage, contrastive learning is utilized to enable the encoder to learn weather-aware representations. The contrastive learning strategy of the WRL stage is based on our proposed multi-queue MoCo, which can tackle the problem of weather class imbalance. The weights of the encoder, the decoder and the queues obtained during the WRL stage will be retained for the CRR stage. In the CRR stage, the encoder is freezed and the main task of this stage is to train a refiner. By treating the normal weather samples as positive examples of contrastive learning, the refiner is guided to pull the adverse weather representations towards normal weather domain. Finally, high-quality density map can be generated to realize precise counting. In terms of the counting head’s structure, we remain consistent with MFA [21], which primarily consists of two transposed convolutional layers.
@@ 3.2 Weather-aware Representation Learning
The model architecture of the WRL stage is illustrated in the upper half of Figure 2. The contrastive learning method is utilized to train the encoders and endow it with the capability to extract weather-aware representations, i.e., images with similar weather conditions correspond to similar representations, whereas those with dissimilar conditions correspond to more distant representations.
Our training structure is similar to that of MoCo [8], which consists of an encoder $\rm E_{Q}$, a momentum-updated encoder $\rm E_{K}$ and a crowd countign head. The encoders $\rm E_{Q}$ and $\rm E_{K}$ extracts representations $\rm R_{Q}$ and $\rm R_{K}$, respectively, from different augmentations $\rm I_{Q}$ and $\rm I_{K}$ of an image. The representations $\rm R_{Q}$ and $\rm R_{K}$ will be further processed by our proposed multi-queue MoCo, which tackles the problem of weather class imbalance and grant the representations with weather-aware nature, to obtain the contrastive loss. The details of multi-queue MoCo will be elaborated in Section 3.3. To ensure that the representation simultaneously contains crowd information, the target of the decoder is set as generating a density map under the supervision of the Bayesian loss [25] according to the representation $\rm R_{Q}$. The total loss of the WRL stage is:
@@ 3.3 Multi-queue MoCo
To tackle the problem, we propose multi-queue MoCo, the architecture of which is illustrated in Figure 3. Similar to most of the contrastive methods, each image undergoes data augmentation multiple times, and the representations originating from the same image as the anchor are treated as positive samples while those from different images are treated as negative samples. This strategy may lead to a situation where samples with same weather conditions from different images are treated as negatives. This seems to be inconsistent with our motivation. But fortunately, research conducted by [36] indicates that contrastive learning has the tolerance to semantically similar negative samples. In light of this, we conducted extensive experiments, which are presented in Section 4.4, and the results indicate that such a strategy outperforms the positive/negative partitioning strategy based on weather labels.
@@ 3.4 Contrastive-learning-guided Representation Refining
The multiple queues obtained during the WRL stage have stored a sufficient number of vectors from various weather conditions. In this paper, we propose to preserve and freeze these queues and employ contrastive learning to guide the refiner to convert the representation. In this stage, all the vectors in the normal-weather sub queue are treated as positive samples and those in other sub queues are treated as negatives. Since there is no longer a need to generate vectors K, the encoder $\rm E_{K}$ and projection head $\rm P_{K}$ are discarded. In order to maintain stable representations, the parameters of encoder $\rm E_{Q}$ are fixed at this stage. The projection head $\rm P_{Q}$ is preserved, fixed, and moved behind the refiner. We denote the vector projected from the refined representation as $\rm Q^{\prime}$. The model structure of the CRR stage is shown in the lower half of Figure 2. The loss function of contrastive learning in this stage is calculated as follows:
@@ 4.1 Experiment Setup
In MQCL, multiple possible architectural configurations can be used as Encoder $\rm E_{Q}$ and $\rm E_{K}$. We employ four distinct models as encoders in our experiments: VGG-16 [29], CSRNet [17], Swin-T [22] and ConvNeXt-T [23]. Except for the removal of the classification head, the architecture of encoders remain identical to the original implementation. The structure of the crowd counting head is consistent with MFA [21], which is an upsampling module composed of two transposed convolutional layers. Two-layer MLPs are employed in the projection heads, the output dimension of which is 2048 and 128. The refiner consists of three ConvNeXt blocks with input and output dimensions of 768, initialized with random parameters.
@@
As mentioned above, each image in a batch undergoes augmentation twice. Specifically, we random crop the image with a size of 256 $\times$ 256, and horizontal flipping is applied for a probability of 50%. In the multi-queue structure, the number of the sub queues is equal to the number of weather classes, and the length of each sub queue is set to 1024. The AdamW optimizer [24] is adopted both in the WRL and the CRR stage, the learning rate is scheduled by a cosine annealing strategy and the initial learning rate is $10^{-4}$. The weight decay is set to $10^{-3}$ and the batch size is $16$. $\lambda_{1}$ and $\lambda_{2}$ in the loss function are both set to 10 and the temperature $\tau$ is set to 0.05.
@@
Following the convention of existing works [17, 19], we adopt Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) as the metrics to evaluate the methods.
@@
Table 1: Quantitative results comparing with the state-of-the-art methods on the JHU-Crowd++ dataset. The best performance is shown in bold and the second best is shown in underlined.
Method Normal Adverse Total
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
SFCN [37] (CVPR 19) 71.4 225.3 122.8 606.3 77.5 297.6
BL [25] (ICCV 19) 66.2 200.6 140.1 675.7 75.0 299.9
LSCCNN [28] (PAMI 20) 103.8 399.2 178.0 744.3 112.7 454.4
CG-DRCN-V [30] (PAMI 20) 74.7 253.4 138.6 654.0 82.3 328.0
CG-DRCN-R [30] (PAMI 20) 64.4 205.9 120.0 580.8 71.0 278.6
UOT [26] (AAAI 21) 53.1 148.2 114.9 610.7 60.5 252.7
GL [34] (CVPR 21) 54.2 159.8 115.9 602.1 61.6 256.5
CLTR [18] (ECCV 22) 52.7 148.1 109.5 568.5 59.5 240.6
MAN [19] (CVPR 22) 46.5 137.9 105.3 478.4 53.4 209.9
AWCC-Net [10] (ICCV 23) 47.6 153.9 87.3 430.1 52.3 207.2
LDFNet [5] (TIP 24) 52.8 156.5 112.4 530.1 59.9 234.8
APGCC [2] (ECCV 24) 48.4 169.3 108.2 478.4 55.5 229.3
Gramformer [20] (AAAI 24) 46.7 170.2 101.2 471.4 53.1 228.1
VGG-16 [29] 68.2 221.4 115.3 632.0 73.8 301.4
MQCL + VGG-16 67.2 209.2 101.4 533.4 71.3 269.3
CSRNet [17] 64.3 223.5 114.8 603.9 70.3 295.8
MQCL + CSRNet 63.1 207.7 99.8 540.2 67.5 269.9
Swin-T [22] 57.8 171.8 103.5 471.7 63.3 229.2
MQCL + Swin-T 50.6 163.8 85.5 458.8 54.8 220.8
ConvNeXt-T [23] (CVPR 22) 52.7 154.9 105.1 561.4 59.0 242.4
MultiModel + ConvNeXt-T 52.9 157.0 108.2 542.2 59.5 238.3
Restore + ConvNeXt-T 52.7 155.9 99.4 533.1 58.3 235.2
MQCL + ConvNeXt-T 48.1 150.0 82.2 433.5 52.1 205.5